OpenStudy (anonymous):

Estimate graphically the local maximum and local minimum of f(x) = 4x^2 + 3x + 2 Local maximum: (-0.4,1.44); local minimum: (0,-0.37) Local maximum: (-0.4,1.44); no local minimum No local maximum; local minimum: (-0.4,1.44) No local maximum; local minimum: (0,-0.37)

OpenStudy (anonymous):

use the vertex form i suppose :)

OpenStudy (anonymous):

I don't knwo it 0:

OpenStudy (primeralph):

Shall I begin?

OpenStudy (anonymous):

You shall

OpenStudy (primeralph):

So me and these chics went dipping and all, and then one of them was like really drunk.

OpenStudy (anonymous):

Omg like no way like forreal ;D

OpenStudy (primeralph):

See, girls always prefer to talk about crazy stuff than to do math. @lornbeach Gotcha!

OpenStudy (anonymous):

I just like to joke around :D

OpenStudy (primeralph):

Of course, me too. Use x = -b/2a

OpenStudy (anonymous):

Can you help me get the function into the vertex form ?

OpenStudy (primeralph):

You don't need that. x = -b/2a = -3/8. a is 4 which is positive, so no local maximum.

OpenStudy (primeralph):

Always a good idea to look at the graph.

OpenStudy (anonymous):

last answer ? @primeralph

OpenStudy (primeralph):

No. That has x = 0. I said x = -3/8.

OpenStudy (anonymous):

which one then.. @primeralph

OpenStudy (primeralph):

It's quite obvious, Dear.

OpenStudy (anonymous):

I guess I'm not seeing it haha

OpenStudy (primeralph):

No maximum. Not D.

OpenStudy (anonymous):

a ?

OpenStudy (primeralph):

Okay, it has no maximum and it is not D. What is it?

OpenStudy (anonymous):

c... ? but c doesn't have 0.375, which is the answer you told me..

OpenStudy (anonymous):

ohhh haha oaky

OpenStudy (primeralph):

Looks like they rounded it.

OpenStudy (primeralph):

The word "estimate" means an approximation.

OpenStudy (anonymous):

Oh okay ! Thanks again (:

OpenStudy (anonymous):

yess